OpenStudy (anonymous):

(x+5)(2x-4)^-1 intervals where the function is increasing or decreasin, inflection points and concavity

OpenStudy (anonymous):

did you take the derivative?

OpenStudy (anonymous):

using quotient rule you will get \[-\frac{7}{(2x-4)^2}\]

OpenStudy (anonymous):

and a little thought will convince you that this thing is always negative, since the denominator is a perfect square.

OpenStudy (anonymous):

therefore the original function is always decreasing. except of course at 2 where it is undefined

OpenStudy (anonymous):

then take the derivative again

OpenStudy (anonymous):

so it does not have a local maximum or minimum?

OpenStudy (anonymous):

this time you get \[\frac{7}{(2x-4)^3}\]

OpenStudy (anonymous):

nope it is always decreasing. try graphing it and you will see

OpenStudy (anonymous):

has vertical asymptote at x = 2, horizontal asymptote at y = 1/2

OpenStudy (anonymous):

you can graph here and see it http://www.wolframalpha.com/input/?i=y%3D%28x%2B5%29%2F%282x-4%29

OpenStudy (anonymous):

always decreasing. we know this because derivative is always negative. so no local max or min

OpenStudy (anonymous):

you can also see from the picture that it is "concave down" on x < 2 and concave up on x > 2

OpenStudy (anonymous):

that is because \[\frac{7}{(2x-4)^3}\] is negative if x < 2 and positive if x > 2

OpenStudy (anonymous):

and the inflection point would be at x=2?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

I don't actually know, there is a discontinuity there, is it still an inflection point?

OpenStudy (anonymous):

as far as i recall "inflection point" means where function changes concavity.

OpenStudy (anonymous):

I think perhaps not, 2 is not in the domain.

OpenStudy (anonymous):

so doesn't need to be defined there. just where it goes from being concave up to concave down or vice versa

OpenStudy (anonymous):

But that is 2, which is not in the domain.

OpenStudy (anonymous):

but i will concede to being wrong for sure

OpenStudy (anonymous):

Normally you want to set the second derivative to 0, a bit difficult here.

OpenStudy (anonymous):

so maybe it is incorrect to say that 2 is an inflection point. maybe it is simply that f is concave down over one interval and concave up over another. i could easily be incorrect.

OpenStudy (anonymous):

ok now i know i am incorrect. http://en.wikipedia.org/wiki/Inflection_point

OpenStudy (anonymous):

Where does it say?

OpenStudy (anonymous):

Ok, I see it now.

OpenStudy (anonymous):

f''(x) should equall zero to have a inflection point?

OpenStudy (anonymous):

Conclusion: function is asymptotic at 2 and there is no point of inflection even though the concavity changes before and after.

OpenStudy (anonymous):

i think the point is that the function should be defined there. it is not necessary for \[f''(x)=0\]

OpenStudy (anonymous):

for example \[f(x)=\sqrt[3]{x}\] has an inflection point at 0 but the derivative is not defined there

OpenStudy (anonymous):

Not every point where the second derivative is zero will be a point of inflection, you need toi understand the overall behavior of the function in order to determine whether it is or not.

OpenStudy (anonymous):

In this case the denominator tells you there is asymptotic behavior at 2 so you should discount a point of inflection there.

OpenStudy (anonymous):

thank you so much for your help